# Decision Machines: An Extension of Decision Trees

### v5

@@ Abstract
Here is a compact representation of binary decision trees. We can explicitly draw the dependencies between prediction and binary tests in decision trees and construct a procedure to guide the input instance from the root to its exit leaf. And we provided a connection between decision trees and error-correcting output codes. Then we built a bridge from tree-based models to attention mechanisms.
@@ 1 Introduction
The conventional induction of decision tree is greedy to recursively partition the training space. The training space is divided into smaller and smaller subsets to minimize the split criteria until a stopping criterion is satisfied for some training data. As a result, it can be represented graphically as a tree. The leaf nodes take the mode of its ground truth set as the prediction value. It is greedy and data-driven. However, the trained decision tree is always implemented as a collection of ‘if-else-then’ sentences. In another word, it is rule-based, which makes it friendly for us to understand and interpret. In contrast, this implementation is unfriendly for computers to execute. The duality, greedy data-driven training and hard-coded rule-based implementation, does not occur in other supervised machine learning methods.
There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [31, 19, 36, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[37].
It is a long history to find diverse representation of decision trees since last century. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a well-known polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 34, 2]. The oblique decision trees are induced from the derivatives of neural models [22].
New representation form would bring or require new optimization and implementation methods. We will show the possibility to reuse the deep learning tool kits to speed up decision trees. Our main contributions include:
  • a provably guaranteed algorithm to represent binary trees in analytic form;
  • a new interpretation of binary decision trees as error correcting output codes;
  • a strong evidence on the possibility to derivate attention mechanisms from decision trees.
@@ 2 Logical Decision Machines
In this section, we will show how to represent the binary decision trees in the language of matrix computation. Our main idea is that the true decision path is the maximum likelihood path over all possible paths. If two input instances arrive at different leaf nodes, they must take different paths and there is at least one different results of the Boolean tests. If they arrive at the same leaf node, they must take the same path, where they must share the same results of all tests. We are looking for an encoding scheme to embed the test results and each decision path into a vector space.
In the rest of this section, we will show how to embed the Boolean tests into a binary space and perform an interleaved traversal of decision trees in the means of matrix computation. And we will provide a specific example and some theoretical insights to the nature of decision trees. Finally, we will prove the guarantee of our algorithm and build a connection between decision trees and error-correcting output codes.
2.1 Construction of Logical Decision Machines
@@
$LS(b,h)=\frac{b\cdot h}{\|b\|_{2}^{2}},$ (3)
where the positive integer $\|b\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $b\in\{1,0,-1\}^{L-1}$. Our similarity(3) is is asymmetric and bounded by $(-1,1]$.
@@
  • If instance $x$ can reach the leaf node $i$ with template vector $b_{i}$, then the similarity $LS(b_{i},h(x))$ is equal to $1$.
  • If instance $x$ can reach the leaf node $i$ with template vector $b_{i}$, then the similarity $LS(b_{i},h(x))$ is the maximum of all possible paths.
@@ 2.2 An Example of the Logical Decision Machine
$Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad h=\operatorname{sgn}(Sx-t)=\begin{pmatrix}+1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},$
$p=\operatorname{diag}(\|B_{1}\|_{1}^{-1},\cdots,\|B_{L}\|_{1}^{-1})Bh=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
@@ 2.4 Provable Guarantee
If $LS(b_{i},h(x))=1$, we can obtain $b_{i},h(x)-b_{i}\right>=0$ by its definition (3) so $h(x)$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $x$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. And the maxima in (4) is unique because $LS(b_{i},h(x))\neq LS(b_{j},h(x))$ when $i\neq j$. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
And we can obtain the convex-concave upper bound on empirical loss proved in [31] and use the non-greedy methods therein to train the logical decision machines.
We call this template matching interpretation of decision trees because it is to select the prediction value based on similarity between template vectors and the result vector of instances.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$. We briefly introduce ECOC in the algorithms(1).
@@
  1. 1.
    Train the binary classifier $f_{s}$ for each column $s=1,\cdots,\ell$.
  2. 2.
    For training classifier $s$, example labeled $y$ is mapped to $M(y,s)$. Omit examples for which $M(y,s)=0$.
  3. 3.
    Given test example $x$, choose the row $y$ of $M$ that is “closest” to binary predictions $(f_{1}(x),\dots,f_{\ell}(x))$ according to some distance (e.g. modified Hamming).
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $B$ corresponds to the error-correcting code matrix $M$. All internal nodes are associated with binary test, which are parameterized by $S$ and $t$. And each binary test corresponds to one dichotomizer. In fact, there is a ternary variant of the error-correcting output codes such as [1]. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $B$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree.
Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaves in decision trees.
@@ 3 Analytical Decision Machines
To show how the probability concentrates on the exit leaf, we reformulate (5) into three equations:
$\begin{split}h&=\bar{B}\operatorname{sgn}(Sx-t),\\ p&=\arg\max_{d\in\Delta}\left<d,h\right>,\\ T(x)&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (6)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{d\mid d\geq 0,d^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $p$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $h\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. The second step in (6) is the key to identify the exit leaf by find the most similar choice. Clearly, the logical decision machine (5) is of hard attention mechanism[35].
@@
$\begin{split}h&=\bar{B}\operatorname{sgn}(Sx-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (7)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\bar{B}$ and $S$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\bar{B}$ and $S$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
@@
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (7) for $i=1,2,\cdots,N$. It is simple to evaluate (8) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T_{1}(x)$ and ${T}_{2}(x)$ of 7 as following
@@ 3.1 Construction of Analytical Decision Machines
The first refinement is to remove the $\operatorname{sgn}$ (1) in (5) or (7). In QuickScorer [25], the exit leaf is identified by the position of the leftmost bit in the product of the false nodes’ bit-vectors, where a bit-vector is assigned to each branch which represents false leaves as $1$ and true leaves as 00. Similarly, we can identify the exit leaf by the position of the rightmost bit in the product of the true nodes’ bit-vectors if we assign a bit-vector to each branch which represents false leaves as 00 and true leaves as $1$. We can identify the exit leaf by the position of the only maximum in the sum of the bit-vectors’ products generated by both algorithms. It is not hard to combine those procedures to remove the $\operatorname{sgn}$ operator in (7) with a well-designed matrix $K$ as described below
$\begin{split}h&=K(Sx-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (9)
where other notation is illustrated in(7). The trouble to remove $\operatorname{sgn}$ in (5) is the probability of zero generated by the affine mapping $Sx-t$.
@@
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; the total $S$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can construct a matrix $M$ to replace $\bar{B}$ in (5) as shown below
@@
Matrix can contain diverse hierarchical information as shown in (5), (8), (9) and (11).
3.2 Attention-Inspired Decision Machines
Great success of the Transformer architecture proposed in [33] benefits from integral combination of its components. In this section, we will lend support to the design choices of Transformer and borrow some ideas to expand the design space of decision machines.
@@ 4 Discussion and Conclusion
We represent the binary decision tree in an analytic form (5). We illustrate this representation with a concrete example and explore some intrinsic properties in theory. Then we prove the equivalence of binary decision trees and two-layer networks with sign activation function 2. And we build bridge between the binary decision tree and the ECOC framework. By explicitly formulating the decision tree, we can obtain the upper bound on empirical loss proved in [31].
We remove the $\operatorname{sgn}$ operator and approximate the $\arg\max$ operator in logical decision machines. Based on the proof of 2 we represent the additive trees in analytic form (8). And we find that the analytic decision machines (10) and (11) are closely related with the attention mechanism.
[…]

### v6

# Decision Machines: Enhanced Decision Trees

@@ Abstract
This paper presents Decision Machines (DMs), an innovative evolution of traditional binary decision trees, which leverages matrix computations to significantly enhance both computational efficiency and interpretability. By explicitly mapping the dependencies between predictions and binary tests within a vector space, DMs offer a streamlined approach to navigating decision paths. We integrate decision trees with kernel methods, ensemble methods and attention mechanisms. The integration of these elements not only bolsters the hierarchical structure of decision trees but also aligns with the computational efficiency of matrix computations. Our work bridges the gap between traditional machine learning algorithms and modern deep learning techniques, providing a novel foundation for further research and application in the field of machine learning.
@@ 1 Introduction
Decision trees have long been a cornerstone in the field of machine learning, revered for their simplicity and interpretability. However, their traditional induction methods, which are inherently greedy and data-driven, often lead to a representation that is less than optimal for computational efficiency. The training process recursively partitions the training space, dividing it into increasingly smaller subsets until a stopping criterion is met, resulting in a tree structure where leaf nodes predict based on the majority of their ground truth set. This approach, while effective, culminates in an implementation that is rule-based and less amenable to computational execution.
In contrast, other supervised learning methods have moved towards end-to-end training schemes, which offer more flexibility and efficiency. The quest for a more computationally friendly representation of decision trees has led to various innovations, including vectorized implementations and attempts to translate decision trees into neural networks. There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [31, 19, 37, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[38]. The oblique decision trees are induced from the derivatives of neural models [22]. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 35, 2]. This paper proposes Decision Machines (DMs), a paradigm that reframes binary decision trees in matrix computation, thereby improving both efficiency and interpretability.
DMs offer a compact representation of binary decision trees, explicitly drawing dependencies between predictions and binary tests. By constructing an arithmetic procedure, we guide input instances from the root to their respective exit leaves, thereby bridging the gap between the intuitive structure of decision trees and the computational efficiency of matrix computations. This formulation not only enhances the interpretability of decision trees but also paves the way for leveraging deep learning toolkits to expedite their execution.
Furthermore, this paper explores the convergence of decision trees with attention mechanisms, a transformative concept in the realm of deep learning. By establishing a connection between tree-based models and attention mechanisms, we open up new avenues for optimizing decision trees and potentially enhancing their predictive power. Our contributions include a provably correct algorithm for representing binary trees analytically, a novel interpretation of binary decision trees as error-correcting output codes, and evidence supporting the derivation of attention mechanisms from decision trees.
@@ 2 Logical Decision Machines
We represent binary decision trees in the language of matrix computation, mapping the true decision path as the maximum likelihood path. The logical decision machines embed Boolean tests into a binary space, enabling an interleaved traversal of decision trees through matrix computation.
2.1 Construction
@@
$LS(k,h)=\frac{k\cdot h}{\|k\|_{2}^{2}},$ (3)
where the positive integer $\|k\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $k\in\{1,0,-1\}^{L-1}$. Our similarity(3) is asymmetric and bounded by $(-1,1]$.
@@
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is equal to $1$.
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is the maximum of all possible paths.
@@
$Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad h=\operatorname{sgn}(Sx-t)=\begin{pmatrix}+1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},\quad p=\bar{B}h=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
@@ 2.2 Provable guarantee
If $LS(b_{i},h(x))=1$, we can obtain $b_{i},h(x)-b_{i}\right>=0$ by the definition (3) so $h(x)$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $x$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. And the maxima in (4) is unique because $LS(b_{i},h(x))\neq LS(b_{j},h(x))$ when $i\neq j$. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
2.3 Consistent interpretation
We can provide some consistent interpretations of the logical decision machines with other machine learning models.
We call the proof 2 the template matching interpretation of decision trees because it is to select the prediction value based on similarity between template vectors and the result vector of instances. We can obtain the convex-concave upper bound on empirical loss proved in [31] and use the non-greedy methods therein to train the logical decision machines.
With the test vectors, we can approximate the prediction value of the $j$th leaf node as a kernel based on random forests estimate in [32]
$\ell_{j}.val=\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}}\frac{y_{i}\delta(\|\vec{t},\vec{t}_{i}\|_{\mathrm{P}})}{N_{j}}$ (6)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise; $\vec{t}$ and $\vec{t}_{i}$ are the test vectors of new instance $\mathbf{x}$ and the $i$-th input instance respectively; the number $N_{j}$ is the total sum of samples falling in the $j$th leaf node; the notation $\|\vec{t},\vec{s}\|_{\mathrm{P}}$ is a distance defined as $\|\mathrm{P}(\vec{t}-\vec{s})\|_{2}$.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$. The encoding stage of ECOC is described in the algorithms(1).
@@ 2.3 Consistent interpretation
1:Input:
  • test example $\mathbf{x}$.
  • the coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$ for class $1,2,\cdots,k$.
  • the pre-trained binary classifiers $f=\{f_{s}\mid s=1,2,\cdots,\ell\}$ for each column of $M$.
2:Output: the predicted label
3:procedure Encoder($\mathbf{x},f$)
4:     Compute the binary predictions of test example $\mathbf{x}$: $p=(f_{1}(x),\dots,f_{\ell}(x))$
5:     for  row $r\in M$ do $\triangleright$ iterate over the rows
6:         $d_{r}\leftarrow\mathrm{dis}(r,p)$
7:     $j\leftarrow$ index of the minimum in $\{d_{1},d_{2},\cdots,d_{\ell}\}$
8:     return $j$
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $B$ corresponds to the error-correcting code matrix $M$. All internal nodes are associated with binary test, which are parameterized by $S$ and $t$. And each binary test corresponds to one dichotomizer. In fact, there is a ternary variant of the error-correcting output codes such as [1]. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $B$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree. Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaf nodes in decision trees.
@@ 3 Analytical Decision Machines
To elucidate the probability concentration on the exit leaf, we have reformulated the process (5) into a more analytical framework, delineating three key equations that capture the essence of our approach.
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(Sx-t),\\ p&=\arg\max_{d\in\Delta}\left<d,h\right>,\\ T(x)&=\sum_{i=1}^{L}p[i]v[i],\end{split}$ (7)
where the notation $\Delta$ is the probability simplex defined as $\Delta=\{d\mid d\geq 0,d^{T}\vec{1}=1\}\subset\mathbb{R}^{L}$; the optima $p$ is in $\{0,1\}^{L}$ where its only non-zero element is in the position of maximum in $h\in\mathbb{R}^{L}$; the integer $L$ is the total number of leaves. Initially, the hidden decision $h$ for an input instance $x$ is calculated through the feature selection matrix $S$ and the threshold vector $t$. Subsequently, the most probable exit leaf is determined by maximizing the dot product within the probability simplex $\Delta$. Finally, the ultimate prediction $T(x)$ is derived as a weighted sum of the values associated with each leaf, reflecting their respective probabilities. Clearly, the logical decision machine (5) is of hard attention mechanism[36].
@@
$\begin{split}h&=\mathrm{\bar{B}}\operatorname{sgn}(\mathrm{S}x-t),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (8)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise. And $\delta(\cdot)$ is activated to identify the exit leaf. Obviously, both $\mathrm{\bar{B}}$ and $\mathrm{S}$ are quite sparse matrices, which may help to understand why sparse or quantized neural networks work. However, the shapes of $\mathrm{\bar{B}}$ and $S$ may restrict the design space of logical decision machines. Additionally, the operators $\operatorname{sgn}(\cdot)$, $\delta(\cdot)$ and $\arg\max(\cdot)$ are non-differentiable, which isolates logical decision machines from deep neural networks.
@@
where $w_{i}\in\mathbb{R}$ and the tree $T_{i}$ is the instances of (8) for $i=1,2,\cdots,N$. It is simple to evaluate (9) in parallel as the same as the random forest [3]. For example, suppose there are two instances $T_{1}(x)$ and ${T}_{2}(x)$ of 8 as following
@@ 3.1 Construction
The first refinement is to remove the $\operatorname{sgn}$ (1) in (5) or (8). The exit leaf is identified by the position of the leftmost bit set to $1$ in the product of the false nodes’ bit-vectors , where are assigned to the internal nodes to represents false leaves as $1$ and true leaves as 00 in QuickScorer [25]. We can combine those procedures and remove the $\operatorname{sgn}$ operator in (8) with well-designed matrices $F$ as described below
$\begin{split}h&=F\operatorname{ReLU}(Wx-b),\\ T(x)&=\sum_{i=1}^{L}\delta(1-h[i])v[i],\end{split}$ (10)
where the operator $\operatorname{ReLU}(Wx-b)$ is defined as $\max(0,Wx-b)$; other notation is illustrated in(8). The trouble to remove $\operatorname{sgn}$ in (5) is the probability of zero generated by the affine mapping $Wx-t$.
@@
where the vector $q=Sx-t$; the real $sim(K_{i},q)_{+}$ is the positive part of the $sim(K_{i},q)$ defined by $\max(0,\left<K_{i},q\right>)$; $S$ is $\sum_{i=1}^{L}sim(K_{i},q)_{+}$; the optional $\alpha$ is a hyperparameter in $(0,1]$. It is a special case of [15].
The third refinement is aimed at realization of the soft decision tree[12]. In [12], each internal node directs instances to its children with probabilities given by a sigmoid gating function. It is equivalent to assign a vector to each branch where left leaves are represented as the sigmoid function $\sigma(\cdot)$; right leaves are represented as $1-\sigma(\cdot)$; and other leaves are represented as $1$. We can derive a matrix $M$ from the template matrix $B$ to replace $\bar{B}$ in (5) as shown below
@@
Matrices in (5), (9), (10) and (12) contain diverse hierarchical information.
3.2 Attention-Inspired models
Great success of the Transformer architecture proposed in [34] benefits from integral combination of its components. In this section, we will lend some support to the design choices of Transformer and borrow some ideas from them to expand the design space of decision machines.
@@ 4 Discussion and Conclusion
In this paper, we have introduced Decision Machines (DMs), a novel extension of binary decision trees, which leverage matrix computation to enhance their computational efficiency and interpretability. By embedding the decision paths and test results into a vector space, DMs offer a more streamlined approach to model prediction. Our exploration into the integration of DMs with attention mechanisms has opened new avenues for leveraging deep learning toolkits to accelerate decision tree operations.